# A Theory of Recommendations

### v2

@@ 7 Conclusion
We have set out to study the role of preference heterogeneity in recommendations, asking questions such as: Under what circumstances do people accept a recommendation? What determines the value of a recommendation system and when is it maximized? Our analysis highlights the importance of disentangling objective information from subjective preferences. We have seen that, depending on the preference distribution, the subjective content can reinforce the positive news of a buy recommendation or reverse it so that a buy recommendation is treated as a bad signal. Conversely, the subjective content can turn a don’t-buy recommendation into good news for some people. The design of a recommendation system, as summarized in the threshold $R$, should thus take into account how it influences the relative objectivity of recommendations.
We find that extreme recommendation systems may be valuable, as they convey objective information even when preferences are heterogeneous. However, when preferences are not distributed symmetrically and controversial products are sufficiently prevalent, with enough consensus on their ranking, the optimal recommendation threshold is set at an intermediate level. This allows receivers to not only draw inferences about the uncontroversial products but also about the controversial ones. For asymmetrically distributed populations, subjective content in the recommendation is thus valuable when controversial products are common and not too controversial.
Throughout the paper, we have analyzed the value of the recommendation system from a consumer perspective. Arguably, online recommendation systems employed on an e-commerce platform are more likely to be geared towards the interests of the platform. Depending on the details of the fee system that it employs, the platform could, for instance, be interested in maximizing the probability of purchases. It is possible to look at such questions in the context of our model, too. A naïve conjecture would be that it is then optimal to set as low a recommendation threshold as possible so that all products except the bad one yield buy recommendations. However, recall that it is not necessarily the case that all receivers accept recommendations. The analysis of this problem is then non-trivial. [^39] In the case of a symmetric population, in which all recommendations are accepted, the problem is simpler, and a sales-maximizing recommendation system indeed sets as low a threshold as possible. Moreover, our results show that this may even be optimal for consumers. This latter part stands in contrast to, for instance, the welfare-reducing inflated ratings in Peitz and Sobolev (forthcoming).
@@ A.1 Proofs and Technical Details for the Benchmark Model
The Effects of Recommendations on Beliefs
To understand how beliefs are updated after a recommendation, it is useful to decompose the difference between beliefs with and without a recommendation into several components. We illustrate this for buy recommendations. Let $\textbf{p}(r)=(p_{H}^{r},p_{1}^{r},p_{2}^{r},p_{L}^{r})$ denote the vector of posterior probabilities following a recommendation $r\in\{B,D\}$. The total effect of the buy recommendation $B$ on the probability vectors is the sum of three parts,
$\textbf{p}(B)-\textbf{q}=(\textbf{p}(B)-\textbf{q}^{\prime\prime})+(\textbf{q}^{\prime\prime}-\textbf{q}^{\prime})+(\textbf{q}^{\prime}-\textbf{q})\text{,}$ (9)
where the three components of the right-hand side will be explained in Steps 1 to 3 below.
Step 1
The first effect of a buy recommendation is that the probability of an objectively bad product falls to zero. We isolate this effect by assuming that the odds ratios between the remaining events remain unchanged so that the effect corresponds to a change from the prior probability q to $\textbf{q}^{\prime}:=(\frac{q_{H}}{1-q_{L}},\frac{q_{1}}{1-q_{L}},\frac{q_{2}}{1-q_{L}},0)$. This effect of the buy recommendation states that all three quality profiles other than $(0,0)$ become more likely at the expense of quality profile $(0,0)$.
Step 2
Next, the buy recommendation means that the objectively good product becomes more likely relative to the controversial products. We isolate this effect by assuming that (i) the probability of $(1,1)$ increases to $p_{H}$ and (ii) the odds ratios between the remaining events remain unchanged so that the effect corresponds to a change from ${\textbf{q}}^{\prime}:=(\frac{q_{H}}{1-q_{L}},\frac{q_{1}}{1-q_{L}},\frac{q_{2}}{1-q_{L}},0)$ to ${\textbf{q}}^{\prime\prime}:=(p_{H},k\frac{q_{1}}{1-q_{L}},k\frac{q_{2}}{1-q_{L}},0)$, where $k=\frac{(1-p_{H})(1-q_{L})}{q_{1}+q_{2}}$ to guarantee that ${\textbf{q}}^{\prime\prime}$ is a probability vector. Clearly, $p_{H}>\frac{q_{H}}{1-q_{L}}$.
Step 3
Finally, the buy recommendation may change the odds between the controversial product versions. We isolate this effect as a change in the probability vector from ${\textbf{q}}^{\prime\prime}:=(p_{H},k\frac{q_{1}}{1-q_{L}},k\frac{q_{2}}{1-q_{L}},0)$ to $\textbf{p}(B)$.
@@
Part (i): Accepting the don’t-buy recommendation is optimal if and only if $U(0)-U(D)\geq 0$ or, equivalently,
$q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)\geq p_{1}^{D}(1/2+i)+p_{2}^{D}(1/2-i)\text{.}$ (10)
Similarly, accepting the buy recommendation is optimal if and only if $U(B)-U(0)\geq 0$ or, equivalently,
$p_{H}^{B}+p_{1}^{B}(1/2+i)+p_{2}^{B}(1/2-i)\geq q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)\text{.}$ (11)
We shall show that conditions (10) and (11) are equivalent, thus proving the claim that receivers either accept all recommendations or never accept recommendations. To do so, we prove the following two steps.
Step 1: (10) is equivalent with
$1-p_{1}^{B}-p_{2}^{D}\geq(1/2+i)[p_{2}^{B}-p_{2}^{D}+p_{1}^{D}-p_{1}^{B}]\text{.}$ (12)
Step 2: (11) is equivalent with (12).
The proofs of both steps will rely on the identities
$q_{s}=\pi^{B}p_{s}^{B}+\pi^{D}p_{s}^{D}=(1-\pi^{D})p_{s}^{B}+\pi^{D}p_{s}^{D}\text{ for }s\in\{H,1,2\}$ (13)
and
$p_{H}^{B}=1-p_{1}^{B}-p_{2}^{B}\text{,}$ (14)
Proof of Step 1: Replacing $q_{H}$, $q_{1}$ and $q_{2}$ in (10) using (13), we obtain
$\left(1-\pi^{D}\right)p_{H}^{B}+\left(\left(1-\pi^{D}\right)p_{1}^{B}+\pi^{D}p_{1}^{D}\right)\left(\frac{1}{2}+i\right)$
$+\left(\left(1-\pi^{D}\right)p_{2}^{B}+\pi^{D}p_{2}^{D}\right)\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\text{.}$
This simplifies to
$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq\left(p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\right)\text{.}$
Using (14) and then rearranging further gives
$1-p_{1}^{B}-p_{2}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\Leftrightarrow$
$1-p_{1}^{B}-p_{2}^{D}-p_{2}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}+i\right)\Leftrightarrow$
$1-p_{1}^{B}-p_{2}^{D}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(-\frac{1}{2}-i\right)\geq\left(p_{1}^{D}-p_{2}^{D}\right)\left(\frac{1}{2}+i\right)\Leftrightarrow$
$1-p_{1}^{B}-p_{2}^{D}\geq(1/2+i)[p_{2}^{B}-p_{2}^{D}+p_{1}^{D}-p_{1}^{B}]\text{.}$
Proof of Step 2: Replacing $q_{H}$ and $q_{1}$ and $q_{2}$ in (11) using (13) gives
$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq$
$\left(1-\pi^{D}\right)p_{H}^{B}+\left(\left(1-\pi^{D}\right)p_{1}^{B}+\pi^{D}p_{1}^{D}\right)\left(\frac{1}{2}+i\right)+\left(\left(1-\pi^{D}\right)p_{2}^{B}+\pi^{D}p_{2}^{D}\right)\left(\frac{1}{2}-i\right)\text{.}$
Rearranging this expression gives
$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-k\right)-p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}-i\right)\geq 0$
Adding $-p_{1}^{B}-p_{2}^{D}+p_{2}^{D}+p_{1}^{B}$ on the left-hand side gives:
$p_{H}^{B}-p_{1}^{B}-p_{2}^{D}+p_{2}^{D}+p_{1}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)-p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}-i\right)\geq 0$
and thus
$p_{H}^{B}-p_{1}^{B}-p_{2}^{D}\geq-p_{2}^{D}-p_{1}^{B}-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\text{.}$
Using (14) gives
$1-p_{1}^{B}-p_{2}^{B}-p_{1}^{B}-p_{2}^{D}$
$\geq-p_{2}^{D}-p_{1}^{B}-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)$
or equivalently
$1-p_{1}^{B}-p_{2}^{B}\geq-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)$
Adding $-p_{2}^{D}+p_{2}^{B}$ on both sides gives:
$1-p_{1}^{B}-p_{2}^{D}-p_{2}^{B}+p_{2}^{B}$
$\geq-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)-p_{2}^{D}+p_{2}^{B}\text{,}$
which is equivalent with (12).
Part (ii): This follows directly from plugging in $\Delta^{r}_{O}$ and $\Delta^{r}_{S}$ into equation (1) for the buy recommendation the analogous equation for a don’t-buy recommendation.
@@
Proof of Lemma 1
@@
Proof of Lemma 2
@@
Proof of Corollary 2
Using (3), applying the $q_{H}$ notation for ease of exposition and imposing Assumption 2 and $q_{1}=q_{2}$ we obtain
$\Delta_{S}^{B}=-\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}.$ (18)
For $a=1$, this term simplifies to $\Delta_{S}^{B}=0$. Hence, it suffices to show that $\Delta_{S}^{B}$ is strictly increasing in $a$. To this end, note that the derivative of $\Delta_{S}^{B}$ with respect to $a$ is given by
$\frac{\partial{\Delta_{S}^{B}}}{\partial{a}}$ $=\frac{Q\left[\log(R)R^{a}\left(2Q(1-R)^{a}+q_{H}\right)+\log(1-R)(1-R)^{a}\left(2Q\left(1-R^{a}\right)+q_{H}\right)\right]}{\left(Q\left(1-R^{a}+(1-R)^{a}\right)+q_{H}\right)^{2}}$
Since $R\in(0,1)$ this expression is negative. Further, given the assumptions, the indifferent type can be expressed as (for $a\neq 1$)
$\tilde{i}$ $=\frac{2Q^{2}\left(1-R^{a}+(1-R)^{a}\right)-2(1-q_{H})q_{H}}{2Q\left(1-R^{a}-(1-R)^{a}\right)}$
$-\frac{Q\left(1-R^{a}+(1-R)^{a}-2q_{H}\left(2-R^{a}+(1-R)^{a}\right)\right)}{2Q\left(1-R^{a}-(1-R)^{a}\right)}$
where
$\frac{\partial{\tilde{i}}}{\partial{Q}}=\frac{Q^{2}\left(1-R^{a}+(1-R)^{a}\right)+(1-q_{H})q_{H}}{Q^{2}\left(1-R^{a}-(1-R)^{a}\right)}$
so that $\tilde{i}$ is increasing in $Q$ when $a>1$ and decreasing in $Q$ when $a<1$.
@@
Using (17) and (18), we obtain
@@
Therefore, $2\Delta_{O}^{B}\geq|\Delta_{S}^{B}|$ so that all types accept a recommendation according to Corollary 1. Thus, Proposition 3 implies
@@
As argued in the text, a mean-preserving spread corresponds to an increase in $\beta$ for $R>1/2$, while for $R<1/2$ it corresponds to a decrease in $\beta$. From the proof of Lemma 2, we can see that the value of the recommendation system is increasing in $\beta$ for $\sigma<1$ and decreasing in $\beta$ for $\sigma>1$. Thus, the statement follows.
Proof of Corollary 3
@@
A.2 Distinct Sender and Receiver Distributions
Proof of Proposition 9

### v3

@@ 7 Conclusion
This paper has examined under which circumstances coarse, non-strategic recommendations can meaningfully guide decision-making in environments with heterogeneous preferences and subjective product quality. Our analysis highlights that even minimal feedback, such as a binary “likes”, can transmit valuable information.
At the heart of our approach is the distinction between the objective and subjective content of a recommendation. While a buy or don’t-buy signal always reveals something about product quality, its usefulness depends on how much of that signal reflects shared preferences versus individual taste. We show that when recommendations are largely objective, all receivers are willing to follow them, either by construction (e.g., through extreme thresholds) or by aggregation (e.g., in symmetric populations).
Systematic preference heterogeneity complicates this picture. In asymmetric populations, where controversial products are common and preferences are aligned in one direction, a recommendation typically contains both objective and subjective content. The same signal—say, a buy recommendation—may reflect a product’s broad appeal, or it may simply reveal that it aligns with the preferences of a particular sender. In such settings, subjectivity introduces friction: Receivers must infer how much of the recommendation reflects shared values rather than idiosyncratic taste, and in some cases, rejecting a recommendation may be the better choice.
However, subjectivity need not be treated solely as a source of noise. When preference alignment is strong and controversial products are sufficiently prevalent, many controversial products are in fact not very controversial in the eyes of the majority. Then, an optimally designed recommendation system enables meaningful inference about such products by preserving some subjective content. In contrast, when this condition fails, it becomes optimal to adopt an extreme threshold that filters out all subjective noise. This stands in contrast to the symmetric population case, where the subjective content of the recommendation system should be suppressed by focusing entirely on extreme thresholds.
Throughout, we contrast the objectives of maximizing user value with maximizing demand. These goals are sometimes aligned, but they often diverge. A platform seeking to drive demand may prefer lenient recommendation thresholds that generate more buy signals, even if this reduces informativeness and harms users.
@@ A.1 Proofs and Technical Details for the Benchmark Model
Posterior Distributions
The receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ after a buy recommendation is
@@
Part (i): Let $\Delta^{B}(i)\equiv U^{B}(i)-U^{0}(i)$ be the gain from following a buy–recommendation and $\Delta^{D}(i)\equiv U^{0}(i)-U^{D}(i)$ the gain from following a don’t-buy–recommendation. A receiver follows a recommendation iff the corresponding gain is non–negative.
For every state $s\in\{H,1,2,L\}$, Bayes’ rule implies
$\pi^{B}(p^{B}_{s}-q_{s})+\pi^{D}(p^{D}_{s}-q_{s})=0.$
Multiplying by the state–contingent payoffs $\{1,1/2+i,1/2-i,0\}$ and summing over the states we obtain
$\pi^{B}\Delta^{B}(i))=\pi^{D}\Delta^{D}(i).$
Because $\pi^{B},\pi^{D}>0$, the differences $\Delta^{D}(i)>0$ and $\Delta^{D}(i)$ have the same sign, so that the receiver either follows both recommendations or rejects both, proving part (i). Part (ii): This follows directly from plugging in $\Delta^{r}_{O}$ and $\Delta^{r}_{S}$ into equation (1) for the buy recommendation the analogous equation for a don’t-buy recommendation.
@@
Proof of Proposition 5
We first restate the probabilities of buy recommendations, their objective content and the value of the recommendation system in terms of $\sigma$, $\beta$ and $Q$:
Lemma A.1.
Suppose Assumption 1 holds. Then,
$\pi^{B}$ $=\frac{(1-2Q)\sigma}{1+\sigma}+2Q\beta;$
Proof.
@@
The following result shows that countervailing effects of $\beta$ on the probability of a buy recommendation and the objective value $\Delta_{O}^{B}$ of accepting it result in a non-trivial effect on $V(\beta)=\pi^{B}\Delta_{O}^{B}$.
Lemma A.2.
Suppose Assumption 1 holds. Then,
  • (i)
  • (ii)
Proof.
@@
Then, using $\beta=F(1/2-R)$, Lemma A.2 directly implies Proposition 5.242424Result (iii) follows because $V(\beta)$ is independent of $\beta$ in this case.
@@
Observe that
$\frac{\partial\tilde{i}}{\partial Q}=\frac{Q^{2}\left(R^{a}-(1-R)^{a}-1\right)+(q_{H}-1)q_{H}}{Q^{2}\left(R^{a}+(1-R)^{a}-1\right)}$ (13)
so that $\tilde{i}$ is monotonically increasing in $Q$ for $a>1$ and monotonically decreasing for $a<1$. Moreover, we have
@@
Therefore, $\tilde{i}$ diverges as $Q\rightarrow 0$. Hence, by continuity and Corollary 1, there exists a cutoff $\bar{Q}$ below which all types accept recommendations. Thus, Proposition 3 implies
@@
We need to show that $D(R)\leq D(0)\forall R>0$. This is equivalent with
$\pi_{B}(R)\hat{F}(R)+(1-\pi_{B}(R))(1-\hat{F}(R))\leq q_{H}+q_{1}+q_{2}$ (14)
To show (14), first note that $q_{H}\geq 1/2$ implies that $\pi_{B}(R)=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)\geq 1/2$. Next, $\pi_{B}(R)\geq 1/2$ is equivalent with
$\pi_{B}(R)\hat{F}(R)+(1-\pi_{B}(R))(1-\hat{F}(R))\leq\pi_{B}(R)$ (15)
From $D(0)=q_{H}+q_{1}+q_{2}$, we further obtain $\pi_{B}(R)\leq D(0)$. Together, this gives (14).
A.2 Results from Section 4.3
Increasing Heterogeneity (Section 4.3.1)
The following result captures the effects of a mean-preserving spread on the value of the recommendation system.
Proposition A.1.
Suppose Assumption 1 holds.
  1. 1.
    If $\sigma<1$, then a mean-preserving spread of the type distribution increases the value of the recommendation system when $R>1/2$ and decreases it when $R<1/2$.
  2. 2.
    If $\sigma>1$, then a mean-preserving spread of the type distribution decreases the value of the recommendation system when $R>1/2$ and increases it when $R<1/2$.
Proof.
As argued in the text, a mean-preserving spread corresponds to an increase in $\beta$ for $R>1/2$, while for $R<1/2$ it corresponds to a decrease in $\beta$. From the proof of Lemma A.2, we can see that the value of the recommendation system is increasing in $\beta$ for $\sigma<1$ and decreasing in $\beta$ for $\sigma>1$. Thus, the statement follows.
To build intuition, suppose $\sigma<1$ and $R>1/2$. As shown in Figure 1, when $R>1/2$, only extreme types issue buy recommendations for controversial products. [^26] For $R<1/2$, only relatively extreme types issue don’t-buy recommendations for controversial products. This follows by reversing the colors in Figure 1 to correspond to $R=1/4$. A mean-preserving spread of $F$ (which corresponds to an increase in $\beta$ given $R$), thus leads to more extreme senders and thus a higher likelihood that buy recommendations stem from controversial products. This raises the probability of buy recommendations while lowering their objective value. The net effect, and whether the system’s value rises or falls, depends on the odds ratio $\sigma$ (see Proposition 5). When $\sigma<1$, good products are rarer than bad ones, making the system more valuable as $\beta$ increases. The benefit of revealing bad products by means of don’t-buy recommendations outweighs the reduced value of buy recommendations. Thus, for high $R$, a more polarized population benefits more from the system, as only bad products result in don’t-buy recommendations. Conversely, for low $R$, a less polarized population centered around the middle ensures that only bad products receive don’t-buy recommendations, thereby yielding a higher value.
Figure 3: Population polarization
[…]
Turning to the impact of preference heterogeneity on demand, the analysis is simpler. Recall that under Assumption 1, all receiver types accept recommendations. Consequently, it follows from Lemma A.2 that demand is increasing in $\beta$. Finally, the same distinction depending on whether $R>1/2$ or $R<1/2$ in terms of the impact of a mean-preserving spread applies as above.
The Prevalence of Controversial Products (Section 4.3.2)
We prove the following result:
Corollary A.1.
Suppose Assumption 1 holds. Then for every $\beta\in(0,1)$, there exist $\underline{\sigma}(\beta)\in(0,1)$ and $\overline{\sigma}(\beta)>1$ such that
  • (i)
    the value of the recommendation system is decreasing in $Q$ on $(\underline{\sigma}(\beta),\overline{\sigma}(\beta))$.
  • (ii)
    For $\sigma<\underline{\sigma}$ and $\sigma>\underline{\sigma}$, the value of the recommendation system is maximized for some $Q^{*}>0$.
Figure 4: Illustration of Corollary 4
(a) Panel A
(b) Panel B
(c) Panel C
Figure 4 illustrates the result. In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists. When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$. For very high or low values of $\sigma$, however, this need not be true. This surprising result reflects the multiplicative structure of $V(\beta)$. Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$. The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B). Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations. However, there are now more buy recommendations stemming from controversial products, as these have become more likely. $\beta$ and $\sigma$ determine which of these effects dominates. When $\beta$ is high, the greater likelihood that controversial products receive buy recommendations increases the total probability of buy recommendations, strengthening the positive effect. When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect. Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C), reflecting the increased probability that a product with a buy recommendation is controversial.  [^27] There is also a positive effect of higher $Q$, reflecting a change in the value of the outside option. The values of $\sigma$ and $\beta$ determine which effect dominates.
Corollary is an immediate implication of the following statement:
Lemma A.3.
Suppose Assumption 1 holds.
  • (i)
  • (ii)
Proof.
@@
A.3 Distinct Sender and Receiver Distributions
We prove the following result:
Proposition A.2.
Consider the model with distinct preference distributions of senders and receivers. Suppose the sender distribution $F$ satisfies Assumption 1. Then, all types accept recommendations. Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
  1. (i)
  2. (ii)
  3. (iii)
Proof.